# Exploiting the diversity of modeling methods to probe systematic biases in strong lensing analyses

### v1

@@ 3.2 Smooth modeling with Sérsic and shapelets
We use a modeling method implemented in the multipurpose Lenstronomy package. We follow the baseline model presented in Sect. 3.1 and which consists of a PEMD plus external shear for the lens. The source is modeled with a Sérsic profile to which are added shapelets components, which is capturing /modeling the azimuthal complexity of the source light distribution. Here, we implement this modeling strategy using version 1.11.3 of the multi-purpose open-source software package Lenstronomy (Birrer & Amara 2018; Birrer et al. 2021). This tool, regularly enhanced with new user-contributed capabilities, provides a large family of lens mass distribution and light profiles. The source is modeled with a Sérsic profile to which is added a set of shapelets components. We refer the reader to Birrer et al. (2015) for a formal description of the shapelets model in the context of lens modeling.
The procedure used in Lenstronomy to derive the posterior distribution on the parameters is sequential. The optimal linear parameters are found through matrix inversion, given values of non-linear parameters. First, a suitable region in non-linear parameter space that minimizes the loss function defined in Eq. 6 is found via a Particle Swarm Optimization algorithm (PSO, Kennedy & Eberhart 2002). Second, the parameters space is sampled using a Monte-Carlo Markov Chain (MCMC). The parameters of the optimal model found previously are randomly perturbed, and used to start the chain. We use the MCMC sampler emcee, which is the most used so far among the Lenstronomy user community (Foreman-Mackey et al. 2013). In this work, as it is also a common practice, the model investigated during the optimization step is a simplified version of the final model, retaining only the main model components that enable to reproduce the largest fraction of the data pixel values. Components that yield small changes of the loss function, such as the source shapelets, are added only during the MCMC sampling. This hierarchy in the significance of model parameters, while not explicitly formalized in the code, is similar to the methodology developed by several automatized lens modeling efforts (e.g., Etherington et al. 2022; Ertl et al. 2023; Tan et al. 2023). We give more technical details regarding this method in Sect. B.1.
@@ 3.3 Adaptive grid source modeling
where $C_{s}$ is some covariance kernel of the source as a function of parameters $\boldsymbol{g}$. In this form, the source parameters can be obtained analytically from $\nabla_{s}L=0$ once the lens parameters $\boldsymbol{\eta}_{\psi}$, regularization strength $\lambda$, and covariance parameters $\boldsymbol{g}$ are given. This approach is referred to as semi-linear inversion. In comparison with forward methods, which may treat more source parameters as non-linear parameters, there are only a few additional parameters that require sampling, $\lambda$ and $\boldsymbol{g}$ (usually corresponding to just one or two parameters), as opposed to tens or hundreds more. The linear source parameters, i.e. pixel brightness values, are obtained analytically.
@@
The regularization of an adaptive grid may be extended to allow for greater fluctuations in surface brightness in the inner regions of the source with high signal-to-noise, while keeping the outer regions of the source relatively smooth. Luminosity-weighted regularization has been explored by Nightingale et al. (2018) in the context of gradient regularization. In this work, we explore luminosity-weighted covariance kernels to achieve a similar effect. To implement this, we define the luminosity-weighted kernel as:
@@
where $\rho$ is a free parameter to be varied, $s_{i,0}$ is an approximate surface brightness of the $i$-th pixel before luminosity weighting, and $s_{\rm max}$ is the maximum surface brightness of all the source pixels. Pixel values $s_{i,0}$ can be taken from the best-fit model of a previous fit altogether, or it can be estimated during each likelihood evaluation by doing an inversion without luminosity weighting first; here, we adopt the latter approach. The advantage of the form given in Eq. 12 is that it ensures that the kernel remains positive-definite, which is critical given the quadratic form of the regularization term. Although more sophisticated forms of the weighting function $W_{i}$ are possible, we only explore the single parameter function given by Eq. 13 in this work.
Here, we use two implementations of the semi-linear inversion technique, the Very Knotty Lenser (VKL, Vernardos & Koopmans 2022) and QLens (Minor et al., in prep.), that can handle different regularization schemes and recipes for constructing the adaptive source-pixel grid. In the VKL modeling run, which we will label Adaptive+Matérn, a Matérn kernel will be used for regularization without supersampling of the image plane; whereas in the QLens modeling runs, an exponential kernel will be used (equivalent to Matérn with $\nu=0.5$), image plane supersampling with $k$-means clustering will be used to construct the adaptive source grid, and models without and with luminosity weighted regularization will be used (Cluster+Exp and Cluster+Exp+Lumweight, respectively). Both implementations use Nested Sampling (as implemented by the MultiNest algorithm, Feroz et al. 2009) to sample the parameter space and converge to the maximum a posteriori solution, in addition to calculating the Bayesian evidence. We outline the specific modeling choices that are considered as three different source modeling techniques in Sects B.2 and B.3.
@@ 3.4 Multi-scale regularization with wavelets
We use the strong lens modeling code Herculens (Galan et al. 2022) that implements the multi-scale regularization strategy described in Eq. 14 using JAX (Bradbury et al. 2018), such that the full model can be pre-compiled and is differentiable. We use the probabilistic programming library NumPyro (Phan et al. 2019; Bingham et al. 2019) to implement prior distributions and constraints on model parameters, and to perform the inference of posterior distributions. Additional technical details regarding the Sparsity+Wavelets models are given in Sect. B.4.
@@ 3.5 Non-parametric Gaussian processes
Finally, we apply a recently introduced source reconstruction method that relies on Gaussian processes and information field theory (IFT Enßlin 2019). The first applications of Gaussian processes for modeling strongly lensed sources are the recent works of Karchev et al. (2022); Rüstig et al. (2024) and Galan et al. (2024), but we describe below the main principles for completeness. In the IFT framework, such Gaussian processes are often referred to as correlated fields, which is the term we use in the remaining.
We use IFT to represent the source light distribution as a two-dimensional non-negative correlated field. Such as field is based on two main components: (1) an analytical parametrization of its power spectrum (mainly its amplitude and slope) which accounts for correlated structures in the source; (2) a discretization onto a regular square grid, with elements following standardized Gaussian distributions (we call the latter an excitation field). More formally, we describe the source field $\boldsymbol{s}$ as
@@
where $\boldsymbol{A}_{0}$ is a zero-mode spectral field generated from the parametrized power spectrum, $\boldsymbol{\xi}$ is the excitation field, and $\odot$ is the point-wise multiplication. The resulting field in harmonic space is then transformed to real space by applying the inverse Fourier transform operator $\boldsymbol{\mathsf{F}}^{-1}$, to which a constant offset $\boldsymbol{\delta}$ is added. Finally, we take the exponential of the resulting field to enforce the positivity of source pixel values, since they represent surface brightness. We note that no extra regularization term is added to the loss function, as Eq. 15 describes a generative model that already incorporates smoothness conditions through its power spectrum.
We use the Python library NIFTy  [^3] https://gitlab.mpcdf.mpg.de/ift/nifty (Selig et al. 2013; Steininger et al. 2019; Arras et al. 2019) to implement the above field model and variational inference samplers to get the joint posterior distribution over the parameter space. As in Galan et al. (2024), we use the JAX interface of NIFTy (nifty.re, Edenhofer et al. 2023) and combine it with Herculens to evaluate the forward model (Eq. 7). For further technical details regarding the Correlated field model, see Sect. B.5.
@@ 3.6 Additional choices left to the modelers
There are some additional choices that are left to the modelers. In particular, they are free to choose if and how they mask out some regions in the imaging data and exclude those from the data likelihood evaluation. Similarly, super-sampling of the coordinates grid when performing ray-tracing evaluations and surface brightness convolutions with the PSF are optional. Modelers are free to run variations of their fiducial models, varying, for example, hyper-parameters, random number generator seeds, or inference algorithms. This may allow for robustness tests, and any eventual marginalization over families of models leading to a joint posterior distribution as the final result. For any further details regarding each modeling method and specific modeling choices, we refer the reader to Appendix B.
@@ 4 Independent simulation of imaging data
We produce a simulated mock system to model with all methods presented above and compare the results. We simulate the mock using MOLET  [^4] https://github.com/gvernard/molet (Vernardos 2022), a simulator code which is independent from any of the codes used to fit the mock. The simulated mock is constructed using a power-law mass profile with external shear and a relatively structured source as expected from most lensed galaxies. The true source used in the simulation is kept hidden (blind) from the modelers, as well as all other input parameters.
For the source, we used an HST image of the galaxy NGC 1084 that was previously used for simulated lenses in the context of the TDLMC (Ding et al. 2021). This source is a local galaxy with a detailed structure that is resolved without any prominent PSF spikes, which could introduce nonphysical features if lensed. We further avoid introducing unphysical features in the resulting lensed source due to edge effects (e.g. sky background) in the cutout that we used by forcing the values in the source pixels to decay exponentially to zero away from the brightness peak.
@@
Simulated lensing data are then created by providing the mass model and source to MOLET, which performs ray-tracing on a high-resolution grid, convolution with the PSF, downsampling, and adding noise to the final mock observation. We use a simulated HST PSF using TinyTim (Krist et al. 2011) based on the WFPC2 instrument with the F814W filter (we do not consider the more recent WFC3 instrument as TinyTim does not support it). Since ray-tracing is performed on a 10 times higher resolution grid compared to the final data, we use a simulated PSF at that resolution for more accurate surface brightness convolutions (although modelers are given a PSF at the data resolution as mentioned in Sect. 3.1). Our settings for the noise correspond to 2200s of exposure in the chosen instrument setup. The angular size of the source cutout is set to 4 arcsec, and we scale its total flux such that it has an apparent (unlensed) AB magnitude of 23.2, which is in the range of observed source galaxies from the SLACS sample (Newton et al. 2011). We show in the top left panel of Fig. 1 the simulated lens image, while the bottom row shows the supersampled and data-resolution PSFs (only the latter is provided to the modelers).
We note that the lensed source galaxy and the data signal-to-noise ratio (S/N) are such that a simple source model is not able to fit the data. We visualize this in the top right panel of Fig. 1, which shows normalized residuals between the data and a model based on a single Sérsic profile for the source. Such residuals are strong evidence for the necessity of more flexible source models as the ones we employ in this work (described in Sect. 3).
In Appendix A we give useful details about a previous version of the mock we attempted to model, for which we detected issues related to the input source light distribution. In a nutshell, the original source did not have an accurate background subtraction and displayed sharp edges (visually unnoticeable after lensing the addition of noise), which lead to biases in the lens models. For this reason, a second mock with different input parameters and source light had to be re-created (and the subsequent modeling re-done).
@@ 5 Standardized comparison framework
Our work relies on several collections of modeling methods and software packages, that we systematically apply on the same data. These codes have been developed following different conventions (e.g., angles, units, profile definitions), are written in different programming languages (e.g., Python, C, C++), and differ in their final modeling products. Therefore, we must ensure that we can both simulate and model strong lensing data in a consistent way, in order to mitigate problems arising from the heterogeneous collection of methods we consider.
We rely on the recently released strong gravitational lensing standard COOLEST (for COde-independent Organized LEnsing STandard) as a framework unifying the different components of our analysis [^5] COOLEST is an open source Python package publicly available at https://github.com/aymgal/COOLEST. We use the released version 0.1.9.. Below we briefly describe COOLEST and its specific features that we use in this work, but refer the reader to Galan et al. (2023) and the online documentation [^6] https://coolest.readthedocs.io for more details.
@@ 6 Results from blind modeling with six methods
During the simulation and subsequent modeling phases, there was minimum amount of information shared between the authors. Keeping this part of the analysis blind ensured unbiased and independent models. One of the authors (M.G.) first simulated the imaging data with a first software (Sect. 4), and the rest of us proceeded with the blind modeling of the data, whose results are presented here. The modeling workload was split between different modelers: L.V.V. (Sérsic+Shapelets models), G.V. (Adaptive+Matérn model), Q.M. (Cluster+Exp and Cluster+Exp+Lumweight models) and A.G. (Sparsity+Wavelets and Correlated Field models). In total, we used four distinct software packages and six different modeling methods. Each modeler was free to perform their analysis to the best of their judgment. Once confident that no significant improvements to the models could be made by further fine-tuning, the modelers converted their results to the COOLEST format (Sect. 5) and submitted them to A.G. (this may have been a single or a marginalization of model instances). At this point, unblinding took place when M.G. also submitted the true model, and the final figures presented in this section were produced. The specific steps necessary to obtain these models, as well as their estimated computation times, are detailed in Appendix B. All models along with the simulated data products are publicly available [^8] https://github.com/aymgal/LensSourceDegeneracy_public. Upon acceptance of the manuscript, analysis and visualization notebooks will also be released..
@@ 6.1 Overall fit to the imaging data
Figure 2: Comparison between all blindly submitted models of the data shown in Fig. 1. The leftmost area gives the labels associated to each modeling method (see also Table 1), as well as their associated color used for the subsequent figures of this paper. First column: image model. Second column: relative difference between the true and modeled convergence maps, with the predicted tangential critical line shown as a black line. Third column: normalized model residuals, with associated reduced chi-squared $\chi_{\nu}^{2}$ indicated in the bottom left. The white areas are outside the likelihood mask chosen by the modelers and are thus excluded during modeling. Fourth column: reconstructed source models, all interpolated onto the same (regular) high-resolution grid to ease visual comparison. The predicted tangential caustic is also indicated as a white line. Last column: reconstructed source models on their original discretization grid, which can be regular or irregular, when applicable. All panels have been generated using COOLEST routines from the standardized storage of each model.
@@
While the Sérsic+Shapelets model captures less small-scale structures in the source compared to other models, the resulting fit in the image plane still achieves noise-level residuals (except maybe for a few pixels in the outskirts of the lensed source). Finer structures in the source like spiral arms and star forming regions are overall are overall better modeled by semi-linear inversion methods, than by the wavelets and the correlated field.

### v2

@@ 3.2 Smooth modeling with Sérsic and shapelets
We used a modeling method implemented in the multipurpose Lenstronomy package. We followed the baseline model presented in Sect. 3.1, and which consists of a PEMD plus external shear for the lens. The source was modeled with a Sérsic profile, to which were added shapelets basis functions, which are capture additional complexity of the source light distribution. We implemented this modeling strategy using version 1.11.3 of the multi-purpose open-source software package Lenstronomy (Birrer & Amara 2018; Birrer et al. 2021). This tool, regularly enhanced with new user-contributed capabilities, provides a large family of lens mass distribution and light profiles. We refer the reader to Birrer et al. (2015) for a formal description of the shapelets model in the context of lens modeling.
The procedure used in Lenstronomy to derive the posterior distribution on the parameters is sequential. The optimal linear parameters are found through matrix inversion, given values of non-linear parameters. First, a suitable region in non-linear parameter space that minimizes the loss function defined in Eq. 6 is found via a Particle Swarm Optimization algorithm (PSO, Kennedy & Eberhart 2002). Second, the parameters space is sampled using a Monte-Carlo Markov Chain (MCMC). The parameters of the optimal model found previously are randomly perturbed, and used to start the chain. We use the MCMC sampler emcee, which is the most used so far among the Lenstronomy user community (Foreman-Mackey et al. 2013). In this work, as it is also a common practice, the model investigated during the optimization step is a simplified version of the final model, retaining only the main model components that enable to reproduce the largest fraction of the data pixel values. Components that yield small changes of the loss function, such as the source shapelets, are added only during the MCMC sampling. This hierarchy in the significance of model parameters, while not explicitly formalized in the code, is similar to the methodology developed by several automatized lens modeling efforts (e.g., Etherington et al. 2022; Ertl et al. 2023; Tan et al. 2024). We give more technical details regarding this method in Sect. B.1.
@@ 3.3 Adaptive grid source modeling
where $C_{s}$ is some covariance kernel of the source as a function of parameters, $\boldsymbol{g}$. In this form, the source parameters can be obtained analytically from $\nabla_{s}L=0$ once the lens parameters, $\boldsymbol{\eta}_{\psi}$, regularization strength, $\lambda$, and covariance parameters, $\boldsymbol{g}$, are given. This approach is referred to as semi-linear inversion. In comparison with forward methods, which may treat more source parameters as non-linear parameters, there are only a few additional parameters that require sampling, $\lambda$ and $\boldsymbol{g}$ (usually corresponding to just one or two parameters). The linear source parameters, in other words the pixel brightness values, are obtained using matrix inversion.
@@
The regularization of an adaptive grid may be extended to allow for greater fluctuations in surface brightness in the inner regions of the source with high S/N, while keeping the outer regions of the source relatively smooth. Luminosity-weighted regularization has been explored by Nightingale et al. (2018) in the context of gradient regularization. In this work, we explored luminosity-weighted covariance kernels to achieve a similar effect. To implement this, we define the luminosity-weighted kernel as:
@@
where $\rho$ is a free parameter to be varied, $s_{i,0}$ is an approximate surface brightness of the $i$-th pixel before luminosity weighting, and $s_{\rm max}$ is the maximum surface brightness of all the source pixels. Pixel values $s_{i,0}$ can be taken from the best-fit model of a previous fit altogether, or it can be estimated during each likelihood evaluation by doing an inversion without luminosity weighting first; here, we adopted the latter approach. The advantage of the form given in Eq. 12 is that it ensures that the kernel remains positive-definite, which is critical given the quadratic form of the regularization term. Although more sophisticated forms of the weighting function $W_{i}$ are possible, we only explored the single parameter function given by Eq. 13 in this work.
Here, we used two implementations of the semi-linear inversion technique, VKL (Vernardos & Koopmans 2022) and QLens (Minor et al., in prep.), which can handle different regularization schemes and recipes for constructing the adaptive source-pixel grid. In the VKL modeling run, which we label Adaptive+Matérn, a Matérn kernel will be used for regularization without supersampling of the image plane; whereas in the QLens modeling runs, an exponential kernel will be used (equivalent to Matérn with $\nu=0.5$), image plane supersampling with $k$-means clustering will be used to construct the adaptive source grid, and models without and with luminosity weighted regularization will be used (Cluster+Exp and Cluster+Exp+Lumweight, respectively). Both implementations use Nested Sampling (as implemented by the MultiNest algorithm, Feroz et al. 2009) to sample the parameter space and converge to the maximum a posteriori solution, in addition to calculating the Bayesian evidence. We outline the specific modeling choices that are considered as three different source modeling techniques in Sects B.2 and B.3.
@@ 3.4 Multi-scale regularization with wavelets
We used the strong lens modeling code Herculens (Galan et al. 2022) that implements the multi-scale regularization strategy described in Eq. 14 using JAX (Bradbury et al. 2018), such that the full model can be pre-compiled and is differentiable. We used the probabilistic programming library NumPyro (Phan et al. 2019; Bingham et al. 2019) to implement prior distributions and constraints on model parameters, and to perform the inference of posterior distributions. Additional technical details regarding the Sparsity+Wavelets models are given in Sect. B.4.
@@ 3.5 Non-parametric Gaussian processes
Finally, we applied a recently introduced source reconstruction method that relies on Gaussian processes and information field theory (IFT Enßlin 2019). The first applications of Gaussian processes for modeling strongly lensed sources are the recent works of Karchev et al. (2022); Rüstig et al. (2024) and Galan et al. (2024), but we describe below the main principles for completeness. In the IFT framework, such Gaussian processes are often referred to as correlated fields, which is the term we use in the remaining.
We use IFT to represent the source light distribution as a two-dimensional non-negative correlated field. Such a field is based on two main components: (1) an analytical parametrization of its power spectrum (mainly its amplitude and slope) that accounts for correlated structures in the source; (2) a discretization onto a regular square grid, with elements following standardized Gaussian distributions (we call the latter an excitation field). More formally, we describe the source field, $\boldsymbol{s}$, as
@@
where $\boldsymbol{A}_{0}$ is a zero-mode spectral field generated from the parametrized power spectrum, $\boldsymbol{\xi}$ is the excitation field, and $\odot$ is the point-wise multiplication. The resulting field in harmonic space is then transformed to real space by applying the inverse Fourier transform operator, $\boldsymbol{\mathsf{F}}^{-1}$, to which a constant offset, $\boldsymbol{\delta}$, is added. Finally, we take the exponential of the resulting field to enforce the positivity of source pixel values, since they represent surface brightness. We note that no extra regularization term is added to the loss function, as Eq. 15 describes a generative model that already incorporates smoothness conditions through its power spectrum.
We use the Python library NIFTy  [^3] https://gitlab.mpcdf.mpg.de/ift/nifty (Selig et al. 2013; Steininger et al. 2019; Arras et al. 2019) to implement the above field model and variational inference samplers to get the joint posterior distribution over the parameter space. As in Galan et al. (2024), we use the JAX interface of NIFTy (nifty.re, Edenhofer et al. 2024) and combine it with Herculens to evaluate the forward model (Eq. 7). For further technical details regarding the Correlated field model, see Sect. B.5.
@@ 3.6 Additional choices left to the modelers
There are some additional choices that are left to the modelers. In particular, they are free to choose if and how they mask out some regions in the imaging data and exclude those from the data likelihood evaluation. Similarly, super-sampling of the coordinates grid when performing ray-tracing evaluations and surface brightness convolutions with the PSF are optional. Modelers are free to run variations in their fiducial models, varying, for example, hyper-parameters, random number generator seeds, or inference algorithms. This may allow for robustness tests, and any eventual marginalization over families of models leading to a joint posterior distribution as the final result.
@@ 4 Independent simulation of imaging data
We produced a simulated mock system to model with all methods presented above and compare the results. We simulated the mock using MOLET  [^4] https://github.com/gvernard/molet (Vernardos 2022), a simulator code that is independent of any of the codes used to fit the mock. The simulated mock is constructed using a power-law mass profile with external shear and a relatively structured source, as is expected from most lensed galaxies. The true source used in the simulation is kept hidden (blind) from the modelers, as well as all other input parameters.
For the source, we used an HST image of the galaxy NGC 1084 that was previously used for simulated lenses in the context of the TDLMC (Ding et al. 2021). This source is a local galaxy with a detailed structure that is resolved without any prominent PSF spikes, which could introduce nonphysical features if lensed. We further avoided introducing unphysical features in the resulting lensed source due to edge effects (e.g., sky background) in the cutout that we used by forcing the values in the source pixels to decay exponentially to zero away from the brightness peak.
@@
The simulated lensing data is then created by providing the mass model and source to MOLET, which performs ray-tracing on a high-resolution grid, convolution with the PSF, downsampling, and adding noise to the final mock observation. We use a simulated HST PSF using TinyTim (Krist et al. 2011) based on the WFPC2 instrument with the F814W filter (we do not consider the more recent WFC3 instrument as TinyTim does not support it). Since ray-tracing is performed on a 10 times higher resolution grid compared to the final data, we use a simulated PSF at that resolution for more accurate surface brightness convolutions (although modelers are given a PSF at the data resolution, as is mentioned in Sect. 3.1). Our settings for the noise correspond to 2200s of exposure in the chosen instrument setup. The angular size of the source cutout is set to 4 arcsec, and we scale its total flux such that it has an apparent (unlensed) AB magnitude of 23.2, which is in the range of observed source galaxies from the Sloan Lens ACS Survey (SLACS) sample (Bolton et al. 2006; Newton et al. 2011). We show in the top left panel of Fig. 1 the simulated lens image, while the bottom row shows the supersampled and data-resolution PSFs (only the latter is provided to the modelers).
We note that the lensed source galaxy and the data S/N are such that a simple source model is not able to fit the data. We visualize this in the top right panel of Fig. 1, which shows normalized residuals between the data and a model based on a single Sérsic profile for the source. Such residuals are strong evidence for the necessity of more flexible source models as the ones we employ in this work (described in Sect. 3).
In Appendix A we give useful details about a previous version of the mock we attempted to model, for which we detected issues related to the input source light distribution. In a nutshell, the original source did not have an accurate background subtraction and displayed sharp edges (visually unnoticeable after lensing the addition of noise), which led to biases in the lens models. For this reason, a second mock with different input parameters and source light had to be re-created (and the subsequent modeling re-done).
@@ 5 Standardized comparison framework
Our work relies on several collections of modeling methods and software packages, that we systematically apply on the same data. These codes have been developed following different conventions (e.g., angles, units, profile definitions), are written in different programming languages (e.g., Python, C++), and differ in their final modeling products. Therefore, we must ensure that we can both simulate and model strong lensing data in a consistent way, in order to mitigate problems arising from the heterogeneous collection of methods we consider.
We used the recently released strong gravitational lensing standard COOLEST (for COde-independent Organized LEnsing STandard) as a framework unifying the different components of our analysis [^5] COOLEST is an open source Python package publicly available at https://github.com/aymgal/COOLEST. We used the released version 0.1.9.. Below we briefly describe COOLEST and its specific features that we used in this work, but refer the reader to Galan et al. (2023) and the online documentation [^6] https://coolest.readthedocs.io for more details.
@@ 6 Results from blind modeling with six methods
After agreeing on the properties of the simulated data (instrumental setup, S/N, type of mass model), one of the authors (M.G.) first simulated the imaging data with a first software (Sect. 4). The rest of the authors proceeded with the blind modeling of the data, whose results are presented here. The modeling workload was split between different modelers: L.V.V. (Sérsic+Shapelets models), G.V. (Adaptive+Matérn model), Q.M. (Cluster+Exp and Cluster+Exp+Lumweight models) and A.G. (Sparsity+Wavelets and Correlated Field models). In total, we used four distinct software packages and six different modeling methods. Each modeler was free to perform their analysis to the best of their judgment. Once confident that no significant improvements to the models could be made by further fine-tuning, the modelers converted their results to the COOLEST format (Sect. 5) and submitted them to A.G. (this may have been a single or a marginalization of model instances). During the simulation and the subsequent modeling phase, there was minimum amount of information shared between the authors. In particular, the shared information between the simulating author and the modelers was restricted to the only data pixels, the corresponding noise map as well as the data-resolution PSF. Afterwards, the modelers did not share their lens modeling results before all authors agree to do so. Keeping this part of the analysis blind ensured unbiased and independent models. Unblinding took place when M.G. also submitted the true model, and the final figures presented in this section were produced. The specific steps necessary to obtain these models, as well as their estimated computation times, are detailed in Appendix B. All models along with the simulated data products are publicly available [^8] https://github.com/aymgal/LensSourceDegeneracy_public. Upon acceptance of the manuscript, analysis and visualization notebooks will also be released..
@@ 6.1 Overall fit to the imaging data
Figure 2: Comparison between all blindly submitted models of the data shown in Fig. 1. The leftmost area gives the labels associated with each modeling method (see also Table 1), as well as their associated color used for the subsequent figures of this paper. First column: Image model. Second column: Relative difference between the true and modeled convergence maps, with the predicted tangential critical line shown as a black line. Third column: Normalized model residuals, with associated reduced chi-squared $\chi_{\nu}^{2}$ indicated in the bottom left. The white areas are outside the likelihood mask chosen by the modelers and are thus excluded during modeling. Fourth column: reconstructed source models, all interpolated onto the same (regular) high-resolution grid to ease visual comparison. The predicted tangential caustic is also indicated as a white line. Last column, first panel: True source as in Fig. 1, to ease comparison with the models. Last column, remaining panels: Reconstructed source models on their original discretization grid, which can be regular or irregular, when applicable (the Sérsic+Shapelets model is not defined on a grid). All panels have been generated using COOLEST routines from the standardized storage of each model.
@@
While the Sérsic+Shapelets model captures less small-scale structures in the source compared to other models, the resulting fit in the image plane still achieves noise-level residuals (except maybe for a few pixels in the outskirts of the lensed source). Finer structures in the source like spiral arms and star forming regions are overall better modeled by semi-linear inversion methods, than by the wavelets and the correlated field.